best suhmesloe for wach question part is used for yeur soore. & nnum form board of length 7 m end mass m-9 k9 i, suspended by twee ropes, if the bansions in the ropes arw 1mgls (on the left) and Bmgt fon the right), what is the location of rm from th·left ropes are OM 1:55 AM TOSHIBA 019 OEN 3 4 5 6 7 8
Answer #1

let the center of mass = x m from left

1mg / 9 + 8mg/ 9 = 9g

1m+ 8m =81

9m = 81

m = 9

tension at left rope= 1mg/ 9 = 9.8 N, tension at right rope= 8x 9,8= 78.4 N

taking moment from left for equilbrium,

9g (x) = 78,4 ( 7)

x = 6.22 m apprx